OpenStudy (anonymous):

-3(-2x+20)+8(x+12)=92

Parth (parthkohli):

I think you might wanna use the distributive property.

OpenStudy (anonymous):

Heya there. Use the distributive property like so: -3*-2x - 3*20 + 8*x +8*12 = 92

OpenStudy (shaik0124):

6x-60+8x+96=92 6x+8x+36=92 14x+36-92=0 14x-56=0 14x=56 x=56/14 x=4

OpenStudy (anonymous):

thank you guys very much!!!

OpenStudy (anonymous):

Np, dude.